### v2

@@ 4.5 Alternative Normalization and Geometric Interpretation
This limiting interpretation provides compelling geometric meaning for the correlation-type co-monotonicity coefficient. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx 1$, the gradients of $y$ and $z$ are nearly parallel at the corresponding point in $M$, indicating that both functions increase in approximately the same direction. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx-1$, the gradients are anti-parallel, so the functions increase in opposite directions. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx 0$, the gradients are approximately orthogonal, indicating that the functions vary independently in different tangent directions. This geometric perspective explains why derivative weighting is natural for smooth functions on manifolds: it captures intrinsic geometric association that is independent of the ambient coordinate system used to represent the data.
Comparison of Normalization Strategies
We now examine how the two normalization strategies differ and when each is preferable for statistical applications. Denote the absolute-value coefficient from Section 4.3 as $\text{cm}_{\text{abs}}(y,z;w)(v)$ and the correlation-type coefficient as $\text{cm}_{\text{cor}}(y,z;w)(v)$. Both have numerators $\sum_{u\in N(v)}w_{e}\Delta_{e}y\cdot\Delta_{e}z$, but different denominators.
The measures answer related but distinct questions about directional concordance. The absolute-value coefficient asks: ”What proportion of directional changes are concordant versus discordant?” It computes a signed sum of weighted products and normalizes by the total absolute magnitude of those products. Each edge contributes to the denominator according to $|w_{e}\Delta_{e}y\cdot\Delta_{e}z|$, which depends on both the weight and the product magnitude. Edges where both functions change substantially receive more weight in the denominator, but the absolute value prevents cancellation between positive and negative products. The resulting measure can be interpreted as the net directional concordance relative to the total amount of joint variation, treating edges democratically in the sense that the normalization accounts for how much each edge contributes to the concordance signal.
The correlation-type coefficient asks: ”How strongly are the directional derivatives correlated?” It normalizes by the geometric mean of squared changes, so the denominator is $\sqrt{\sum w_{e}(\Delta_{e}y)^{2}}\sqrt{\sum w_{e}(\Delta_{e}z)^{2}}$. This treats the functions separately before combining: we measure the total variation of each function independently, then correlate their directional changes. Edges where one or both functions vary substantially contribute more to their respective denominator terms, making the measure sensitive to the magnitude profile of changes. This aligns with classical Pearson correlation, which measures linear association weighted by the variances of the two variables.
By the Cauchy-Schwarz inequality applied to the sequences $(a_{u}=\sqrt{w_{e}}\Delta_{e}y)_{u\in N(v)}$ and $(b_{u}=\sqrt{w_{e}}\Delta_{e}z)_{u\in N(v)}$, we have
$\sum_{u\in N(v)}w_{e}|\Delta_{e}y\cdot\Delta_{e}z|\leq\sqrt{\sum_{u\in N(v)}w_{e}(\Delta_{e}y)^{2}}\sqrt{\sum_{u\in N(v)}w_{e}(\Delta_{e}z)^{2}},$ (32)
with equality if and only if $|\Delta_{e}y|$ and $|\Delta_{e}z|$ are proportional across all edges: $|\Delta_{e}z|=k|\Delta_{e}y|$ for some constant $k\geq 0$ independent of $u$. When the numerator is positive (indicating net concordance), dividing both sides by the numerator and taking reciprocals shows that
$|\text{cm}_{\text{cor}}(y,z;w)(v)|\leq|\text{cm}_{\text{abs}}(y,z;w)(v)|,$ (33)
with equality when changes are proportional in magnitude across edges. The absolute-value coefficient generally has larger magnitude because its denominator is smaller (or equal), reflecting that it requires less total variation to achieve a given level of net concordance.
The measures coincide in specific situations that reveal when normalization choice is inconsequential. Perfect proportionality $\Delta_{e}z=k\Delta_{e}y$ for constant $k$ across all edges incident to $v$ produces $\text{cm}_{\text{abs}}=\text{cm}_{\text{cor}}=\text{sign}(k)$, since both denominators equal $|k|$ times the same quantity. When all edge weights are equal and all magnitudes $|\Delta_{e}y|$ and $|\Delta_{e}z|$ are approximately constant across edges, the two normalizations differ only by a constant factor that cancels in the ratio. These scenarios represent homogeneous neighborhoods where all edges provide comparable information about directional concordance.
The measures diverge most strongly when magnitude heterogeneity is substantial. Consider a vertex $v$ with two neighbors $u_{1}$ and $u_{2}$, where $\Delta_{e_{1}}y=0.1$, $\Delta_{e_{1}}z=0.1$, while $\Delta_{e_{2}}y=0.1$, $\Delta_{e_{2}}z=5$, with unit edge weights. Both edges show concordance (positive products), but the second edge exhibits much larger magnitude change in $z$. The absolute-value coefficient computes
$\text{cm}_{\text{abs}}=\frac{0.01+0.5}{0.01+0.5}=1,$ (34)
indicating perfect concordance since both edges agree in direction. The correlation-type coefficient gives
$\text{cm}_{\text{cor}}=\frac{0.01+0.5}{\sqrt{0.01+0.01}\sqrt{0.01+25}}=\frac{0.51}{\sqrt{0.02}\sqrt{25.01}}\approx 0.228,$ (35)
a much smaller value reflecting that the large change in $z$ on the second edge is not matched by a correspondingly large change in $y$, indicating imperfect correlation of magnitudes despite consistent direction. The absolute-value measure emphasizes directional agreement, while the correlation measure penalizes magnitude mismatch.
For statistical applications, the choice between normalizations depends on the inferential goals and data characteristics. We recommend using $\text{cm}_{\text{abs}}$ when robustness to occasional extreme edges is desired. In noisy data where individual edges may exhibit large but unreliable changes due to measurement error or sampling artifacts, the absolute-value normalization prevents such edges from dominating the association measure. The frequency-based interpretation (proportion of concordant changes) remains meaningful even when individual change magnitudes are uncertain. This normalization is also preferable when the scientific question focuses on consistency of directional relationships rather than strength of linear association.
The correlation-type $\text{cm}_{\text{cor}}$ is preferable when edges represent measurements of varying precision and larger changes carry more information. In applications where the magnitude of change reflects confidence or biological significance, weighting by squared magnitudes appropriately up-weights reliable signals. The geometric interpretation as cosine of gradient angles provides theoretical grounding and connects to classical differential geometry. When downstream analyses assume correlation-like structure (such as principal component analysis or factor models applied to vertex-level coefficients), using $\text{cm}_{\text{cor}}$ ensures compatibility with those methods.
In practice, computing both measures and examining their differences provides diagnostic information about local association structure. Vertices where $|\text{cm}_{\text{abs}}|\gg|\text{cm}_{\text{cor}}|$ indicate regions where directional concordance is consistent but magnitude matching is poor, suggesting that the association may be monotonic but non-linear. Vertices where the two measures are comparable suggest homogeneous neighborhoods with proportional changes, indicating robust local association. Vertices where both measures are small indicate weak or inconsistent relationships regardless of normalization. This dual perspective enriches the analysis by revealing not just whether functions co-vary but how the pattern of co-variation manifests across neighborhood edges.
The availability of two complementary measures also enables sensitivity analysis for downstream inference. If regional clustering or hypothesis testing produces similar results using either $\text{cm}_{\text{abs}}$ or $\text{cm}_{\text{cor}}$, this provides evidence that the findings are robust to normalization choice and likely reflect genuine biological or scientific patterns rather than artifacts of measurement scale. Conversely, if results differ substantially, this flags vertices or regions where conclusions depend critically on whether we prioritize directional consistency versus magnitude correlation, prompting closer examination of the underlying data and scientific interpretation.
@@ 4.6 Matrix Extension for Multivariate Analysis
Computing $\text{CM}(y,Z)$ naively by $m$ independent calls to the pairwise co-monotonicity function would redundantly recompute edge differences $\Delta_{e}y$ for each feature. We optimize by precomputing these $y$-dependent quantities once, then reusing them across all features. The algorithm maintains edge weights $w_{e}$ and differences $\Delta_{e}y$ in memory, iterates over features $z_{j}$, and for each feature computes $\Delta_{e}z_{j}$ and accumulates weighted products. This reduces computational cost from $O(m\cdot|E|)$ operations performed independently to $O(|E|+m\cdot|E|)$ with shared overhead, approximately halving the runtime for large $m$.
@@
In applications, we augment $\text{CM}(y,Z)$ with selected columns from $\text{CM}(Z,Z)$ to form an extended association profile. For a vertex $v$, this profile includes both $\text{cm}(y,z_{j})(v)$ for all features $j$ and $\text{cm}(z_{j},z_{k})(v)$ for pairs $(j,k)$ of interest. These profiles embed vertices into an association space where proximity reflects similarity in how outcome and features relate locally. Vertices with similar profiles form co-monotonicity cells, which we will explore in subsequent sections as a basis for geometric multiple testing and biclustering.

### v3

@@ 4.3 Geometric Interpretation via Gradient Alignment
This limiting interpretation provides compelling geometric meaning for the correlation-type co-monotonicity coefficient. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx 1$, the gradients of $y$ and $z$ are nearly parallel at the corresponding point in $M$, indicating that both functions increase in approximately the same direction. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx-1$, the gradients are anti-parallel, so the functions increase in opposite directions. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx 0$, the gradients are approximately orthogonal, indicating that the functions vary independently in different tangent directions.
This geometric perspective explains why derivative weighting is natural for smooth functions on manifolds: it captures intrinsic geometric association that is independent of the ambient coordinate system used to represent the data. The normalization by edge length converts raw differences into discrete approximations of directional derivatives, which are the natural geometric quantities for characterizing function variation on Riemannian manifolds.
In finite samples with discrete neighborhoods, this limiting result suggests why the correlation-type formulation performs well in practice. The asymptotic formula assumes the neighborhood becomes dense in the tangent space, providing approximately uniform directional coverage. With finite $k$, the discrete neighborhood necessarily undersamples certain tangent directions, but the correlation-type normalization remains stable provided both functions exhibit sufficient variation across incident edges. The geometric mean in the denominator ensures the coefficient scales appropriately even when gradient magnitudes vary spatially, making the measure robust to heterogeneous signal patterns common in high-dimensional biological data.
4.4 Edge Weighting Schemes
The correlation-type co-monotonicity coefficient defined in Section 4.2 allows flexible edge weighting through the parameters $w_{e}\geq 0$. The choice of weights determines which aspects of directional concordance the measure emphasizes, with different schemes appropriate for different data structures and inferential goals. We examine three canonical weighting schemes that arise naturally in geometric data analysis.
The simplest choice assigns uniform weight to all edges: $w_{e}=1$ for all $e\in E$. This unit weighting scheme treats edges equally regardless of their length, emphasizing the combinatorial structure of the graph. The resulting coefficient
$\text{cm}_{\text{cor}}(y,z;1)(v)=\frac{\sum_{u\in N(v)}\Delta_{e}y\cdot\Delta_{e}z}{\sqrt{\sum_{u\in N(v)}(\Delta_{e}y)^{2}}\sqrt{\sum_{u\in N(v)}(\Delta_{e}z)^{2}}}$ (29)
directly correlates the raw function differences across edges without geometric normalization. Unit weighting is appropriate when edge lengths are roughly comparable, as in regular lattices or uniformly sampled grids, or when the graph represents abstract relationships rather than geometric proximity. For graphs constructed from irregular spatial samples where edge lengths vary substantially, unit weighting may overweight contributions from long edges that span large distances in the underlying space, obscuring local patterns of directional concordance.
To account for geometric scaling, we employ derivative weighting as introduced in Section 4.2. Setting $w_{e}=1/\ell_{e}$ where $\ell_{e}$ denotes edge length yields
$\text{cm}_{\text{cor},\partial}(y,z)(v)=\frac{\sum_{u\in N(v)}\frac{\Delta_{e}y}{\ell_{e}}\cdot\frac{\Delta_{e}z}{\ell_{e}}}{\sqrt{\sum_{u\in N(v)}\left(\frac{\Delta_{e}y}{\ell_{e}}\right)^{2}}\sqrt{\sum_{u\in N(v)}\left(\frac{\Delta_{e}z}{\ell_{e}}\right)^{2}}},$ (30)
which correlates discrete directional derivatives rather than raw differences. This normalization by edge length has two beneficial effects. First, it prevents long edges from dominating the coefficient: a large absolute change $\Delta_{e}y$ across a long edge $\ell_{e}$ corresponds to a moderate derivative $\Delta_{e}y/\ell_{e}$, giving it appropriate weight relative to shorter edges. Second, as established in Section 4.3, derivative weighting ensures that the coefficient approximates the cosine of the angle between gradient vectors on smooth manifolds, providing intrinsic geometric meaning independent of coordinate representation. Derivative weighting is appropriate when functions represent continuous quantities sampled at irregular positions and we wish to assess co-variation in the underlying continuous fields rather than artifacts of the sampling pattern.
A third weighting scheme arises naturally from spectral graph theory and diffusion processes on graphs. In the graph Laplacian framework, edge weights often represent conductances $c_{e}$ that govern the rate of diffusion between adjacent vertices. Using $w_{e}=c_{e}$ yields a coefficient that emphasizes edges with high conductance, reflecting the effective connectivity structure rather than purely geometric proximity. For graphs derived from kernel-weighted similarity matrices where $c_{e}=\exp(-d_{e}^{2}/\sigma^{2})$ with $d_{e}$ denoting distance and $\sigma$ controlling bandwidth, conductance weighting downweights distant connections while preserving strong local relationships. This scheme proves useful when the graph encodes probabilistic or functional relationships rather than strict geometric structure, as in biological networks where interaction strength varies independently of physical distance.
The choice among these weighting schemes depends on the data structure and the inferential question. Derivative weighting is generally preferred for spatially embedded data where geometric scaling matters and functions are approximately continuous. Unit weighting suffices for combinatorial graphs or when edge lengths are comparable. Conductance weighting addresses situations where edge importance varies beyond geometric considerations, capturing the effective strength of relationships encoded by the graph structure. In all cases, the correlation-type normalization ensures coefficients remain bounded in [1,1][-1,1] and retain interpretability as measures of directional concordance scaled by the geometric mean of function variations.
4.5 Scale Artifacts and Geometric Smoothing
The unit and derivative weighting schemes measure directional concordance at different geometric scales, which can produce substantial discrepancies when graph edge lengths vary heterogeneously. We demonstrate this phenomenon through an empirical analysis of spontaneous preterm birth associations in vaginal microbiome data (Figure 4), then introduce geometric smoothing as a principled approach for scale-consistent inference.
Consider a k-nearest neighbor graph constructed from compositional data, where edge lengths $\ell_{e}$ represent distances between samples in a high-dimensional feature space. The distribution of edge lengths typically exhibits considerable heterogeneity: most edges connect samples at moderate distances (say, $0.05<\ell_{e}<0.15$), but some edges span very short distances ($\ell_{e}<0.01$) between nearly identical samples arising from repeated measurements, high sampling density in certain regions, or genuine biological clustering. Under derivative weighting with $w_{e}=1/\ell_{e}$, these very short edges receive extreme weights that can exceed typical edge weights by factors of $10^{3}$ to $10^{5}$.
We assessed the practical impact of scale heterogeneity by computing both unit-weighted and derivative-weighted correlation coefficients between sPTB prevalence and 106 phylotype abundances across 2,117 vertices in a vaginal microbiome graph, yielding 224,202 vertex-phylotype pairs. Figure 4A displays the relationship between these two measures for all pairs. While most pairs cluster near the diagonal, indicating agreement between weighting schemes, substantial discrepancies appear at 964 pairs (0.43% of total, highlighted in red) where $|\text{cm}_{\text{cor}}(y,z;1)(v)-\text{cm}_{\text{cor},\partial}(y,z)(v)|>0.75$. The scatter reveals that unit and derivative weighting can yield markedly different assessments of the same biological association. One extreme case displayed opposite signs: unit weighting yielded 0.618-0.618 while derivative weighting yielded +0.906+0.906 for the association between sPTB and Corynebacterium imitans at a specific vertex.
Detailed examination of this extreme case revealed the mechanism underlying such discrepancies. The vertex had 37 neighbors, but five edges with lengths below 0.005 received derivative weights exceeding 200-fold typical values, dominating the calculation and yielding opposite-sign concordance from the broader neighborhood pattern. The unit-weighted coefficient averaged uniformly across all 37 edges, yielding a negative association reflecting the majority pattern. The derivative-weighted coefficient was driven almost entirely by the five extremely short edges, which exhibited positive concordance while the remaining 32 edges showed negative concordance.
[Uncaptioned image][Uncaptioned image]
Figure 4: Scale artifacts in co-monotonicity coefficients are eliminated by geometric smoothing. (A) Raw correlation-type coefficients computed with derivative versus unit weighting for sPTB prevalence and 106 phylotypes across 224,402 vertex-phylotype pairs. Red points indicate pairs with large discrepancies ($|\text{difference}|>0.75$). (B) The same coefficients after geometric smoothing via graph Laplacian filtering. Smoothing eliminates scale-dependent discrepancies, demonstrating that both weighting schemes recover nearly identical association structure at consistent geometric scale.
This behavior reflects fundamentally different notions of local association. Unit weighting measures concordance across the combinatorial neighborhood, treating all connected vertices equally regardless of their geometric separation. Derivative weighting measures concordance in the geometric tangent space, where very short edges approximate infinitesimal neighborhoods and receive proportionally large weight according to the distance gradient. When micro-scale patterns at edges shorter than 0.01 differ from meso-scale patterns at edges between 0.05 and 0.15, the two measures capture genuinely distinct geometric phenomena operating at different scales. The derivative-weighted coefficient becomes sensitive to association structure at arbitrarily fine scales determined by the shortest edges in the graph, which may reflect sampling artifacts rather than meaningful biological variation. Conversely, the unit-weighted coefficient treats all scales equally within the combinatorial neighborhood, potentially obscuring geometric structure encoded in edge lengths. Without additional constraints, neither measure uniquely captures the associations of scientific interest.
These findings demonstrate that raw co-monotonicity coefficients exhibit scale-dependent artifacts that obscure biological signal. We address this scale heterogeneity through geometric smoothing of the computed coefficient matrices. The graph Laplacian low-pass filter employed for smoothing response functions (Section 3) naturally extends to smoothing association maps. For a co-monotonicity matrix $\text{CM}(y,Z)$ with dimensions $n\times m$, we apply the smoothing operation to each column independently:
$\text{CM}_{\text{smooth}}(y,Z)=[\text{smooth}(\text{CM}(y,Z)[:,j])]_{j=1,\ldots,m},$ (31)
where the smoothing uses the same regularization parameter $\lambda$ selected for the response $y$ via generalized cross-validation or similar criteria. This ensures that associations are analyzed at the same geometric scale as the response variation itself, maintaining consistency throughout the analysis pipeline.
Figure 4B displays the result of applying this smoothing procedure to both unit-weighted and derivative-weighted coefficients. The effect is substantial: the mean absolute difference between the two measures drops from 0.089 to 0.029, the 95th percentile decreases from 0.329 to 0.095, and all 964 pairs with raw differences exceeding 0.75 reduce to disagreements below this threshold after smoothing. The smoothed coefficients align almost perfectly along the diagonal, with correlation exceeding 0.999 between the two measures. The extreme case of opposite signs resolved to agreement: both smoothed coefficients yielded +0.82+0.82, confirming a positive association at the biologically relevant scale while filtering out the micro-scale artifact from very short edges. This near-perfect concordance after smoothing demonstrates that unit and derivative weighting recover the same underlying association structure once analyzed at a consistent geometric scale.
Figure 5: Raw unit-weighted coefficients do not match smoothed derivative-weighted coefficients, demonstrating that smoothing fundamentally transforms both weighting schemes. Comparison of raw unit-weighted co-monotonicity coefficients versus smoothed derivative-weighted coefficients for sPTB prevalence and 106 phylotypes across 224,402 vertex-phylotype pairs. The wide scatter (Gini mean difference: 0.322) contrasts with the tight agreement between smoothed unit-weighted and smoothed derivative-weighted coefficients (Figure 4B), confirming that geometric smoothing modifies association structure rather than merely adjusting one scheme to match the other.
Importantly, both unit-weighted and derivative-weighted coefficients are substantially modified by the smoothing process. Figure 5 compares raw unit-weighted coefficients with smoothed derivative-weighted coefficients, revealing persistent discrepancies throughout the distribution. While the mean difference is near zero (0.007-0.007), substantial scatter remains with the 5th and 95th percentiles at 0.490-0.490 and $0.453$ respectively, range from 1.69-1.69 to $1.57$, and Gini mean difference of $0.322$. This wide scatter contrasts sharply with the tight agreement between smoothed unit-weighted and smoothed derivative-weighted coefficients (Figure 4B), confirming that geometric smoothing substantively modifies the association structure for both weighting schemes rather than simply adjusting one to match the raw values of the other.
This empirical validation reveals that the apparent conflict between unit and derivative weighting reflects scale artifacts rather than fundamental methodological differences. The biologically meaningful associations exist at the geometric scale determined by the graph Laplacian regularization, which balances signal fidelity against noise suppression. Micro-scale variations captured by extreme derivative weights on very short edges represent either sampling artifacts such as repeated measurements and density fluctuations, or true but scientifically irrelevant local fluctuations. The smoothing operation removes precisely these high-frequency components, leaving the large-scale association structure that both normalizations agree upon. This convergence at consistent geometric scales indicates that the choice of weighting scheme becomes largely immaterial after appropriate smoothing. Both smoothed unit-weighted and smoothed derivative-weighted coefficients provide equally defensible measures of association structure. We adopt smoothed derivative-weighted coefficients for subsequent analyses, as derivative weighting naturally adapts to local manifold geometry, but we acknowledge this choice is pragmatic rather than definitive.
@@ 4.6 Matrix Extension for Multivariate Analysis
Computing $\text{CM}(y,Z)$ naively by $m$ independent calls to the pairwise co-monotonicity function would redundantly recompute edge differences $\Delta_{e}y$ for each feature. We optimize by precomputing these $y$-dependent quantities once, then reusing them across all features. The algorithm maintains edge weights $w_{e}$ and differences $\Delta_{e}y$ in memory, iterates over features $z_{j}$, and for each feature computes $\Delta_{e}z_{j}$ and accumulates weighted products. This reduces computational cost from $O(m\cdot|E|)$ operations with independent overhead to $O(|E|+m\cdot|E|)$ with shared preprocessing, yielding substantial savings when $m$ is large.
@@
In applications, we augment $\text{CM}(y,Z)$ with selected columns from $\text{CM}(Z,Z)$ to form an extended association profile. For a vertex $v$, this profile includes both $\text{cm}(y,z_{j})(v)$ for all features $j$ and $\text{cm}(z_{j},z_{k})(v)$ for pairs $(j,k)$ of interest. These profiles embed vertices into an association space where proximity reflects similarity in how outcome and features relate locally. Vertices with similar profiles form co-monotonicity cells, which we will explore in Section 5 as a basis for geometric multiple testing and biclustering.
4.7 Alternative Normalizations
The correlation-type co-monotonicity coefficient developed in Sections 4.2-4.4 employs Pearson-style normalization, dividing by the geometric mean of squared variations. This normalization provides stable estimates and interpretable coefficients bounded in [1,1][-1,1] for most applications. However, specific data characteristics or inferential goals may suggest alternative normalization schemes. We examine three such alternatives, providing concrete examples that illustrate when each approach succeeds or fails.
We begin by considering absolute value normalization, which divides the sum of weighted products by the sum of weighted absolute products. For edge weights $w_{e}$, this yields
$\text{cm}_{\text{abs}}(y,z;w)(v)=\frac{\sum_{u\in N(v)}w_{e}\Delta_{e}y\cdot\Delta_{e}z}{\sum_{u\in N(v)}w_{e}|\Delta_{e}y\cdot\Delta_{e}z|}.$ (34)
This normalization has intuitive interpretation: it measures the balance between concordant edges (positive products) and discordant edges (negative products), weighted by the magnitude of their products. When all edges show agreement, the numerator equals the denominator and $\text{cm}_{\text{abs}}=1$. When disagreements exactly balance agreements in product-weighted terms, the coefficient vanishes.
The critical weakness of absolute value normalization emerges when one function varies sparsely. Consider a vertex $v$ with 30 neighbors where the outcome $y$ exhibits variation of magnitude 0.1 across all edges, while feature $z$ shows meaningful variation (magnitude 0.1) on only one edge and remains essentially constant (changes <0.01<0.01) on the remaining 29 edges. The single edge with signal contributes product $\pm 0.01$ to the numerator. The 29 near-constant edges contribute products near zero. The denominator sums the absolute values of these products, yielding approximately $0.01+29\times 0.001=0.039$. The resulting coefficient is $\pm 0.01/0.039\approx\pm 0.26$, suggesting moderate association despite signal appearing on only $1/30\approx 3\%$ of edges. As the non-signal edges become exactly constant, the coefficient approaches $\pm 1$, indicating perfect association based on a single edge. This instability arises because the denominator can become arbitrarily small when one function is nearly constant, allowing sparse signal to dominate the measure.
Sign-based normalization addresses magnitude sensitivity by considering only directional information. We replace the weighted sums with counts of edges showing agreement versus disagreement:
$\text{cm}_{\text{sign}}(y,z)(v)=\frac{\sum_{u\in N(v)}\text{sign}(\Delta_{e}y\cdot\Delta_{e}z)}{|N(v)|},$ (35)
where $\text{sign}(x)=1$ if $x>0$, $\text{sign}(x)=-1$ if $x<0$, and $\text{sign}(0)=0$. This coefficient measures the proportion of edges where functions agree in direction minus the proportion where they disagree, treating all edges equally regardless of change magnitude.
Sign-based normalization discards potentially valuable information about the strength of associations. Consider two scenarios at a vertex with 20 neighbors. In the first scenario, $y$ and $z$ both change by 0.5 units in the same direction on 11 edges and in opposite directions on 9 edges, yielding substantial co-variation with correlation 0.7. In the second scenario, $y$ changes by 0.5 units on all edges while $z$ shows tiny changes of 0.01 units in directions that happen to align with $y$ on 11 edges and oppose $y$ on 9 edges, exhibiting negligible true co-variation. The sign-based coefficient produces (119)/20=0.1(11-9)/20=0.1 in both cases, treating large coordinated changes identically to small random fluctuations. The magnitude information that distinguishes biologically meaningful co-variation from numerical noise is completely lost. Moreover, sign-based coefficients remain vulnerable to the sparse signal problem: when $z$ varies on only one edge, that single edge’s sign determines the coefficient, yielding values of $\pm 1/|N(v)|$ that overstate association strength relative to the prevalence of signal.
These limitations of absolute value and sign-based normalizations arise from their treatment of edges with weak or absent signal. The absolute value denominator becomes small when products are small, amplifying sparse signals. The sign-based approach counts edges with arbitrarily weak variation equally with edges showing strong variation, obscuring signal strength. In practice, neither normalization provides reliable inference for heterogeneous data where features exhibit spatially varying signal intensity. We do not recommend their use in applications requiring robust, interpretable association measures.
For scenarios involving sparse or binary features where explicit signal filtering is scientifically justified, we introduce proportion-based co-monotonicity with thresholds. This measure explicitly filters edges by signal strength before aggregating directional agreements. Given threshold parameters $\tau_{y},\tau_{z}>0$, we assign each edge $e=[v,u]$ a directional agreement score:
$s_{e}=\begin{cases}+1&\text{if }|\Delta_{e}y|>\tau_{y},|\Delta_{e}z|>\tau_{z},\text{ and }\Delta_{e}y\cdot\Delta_{e}z>0\\ -1&\text{if }|\Delta_{e}y|>\tau_{y},|\Delta_{e}z|>\tau_{z},\text{ and }\Delta_{e}y\cdot\Delta_{e}z<0\\ 0&\text{otherwise.}\end{cases}$ (36)
The proportion-based coefficient at vertex $v$ is then
$\text{cm}_{\text{prop}}(y,z;\tau_{y},\tau_{z})(v)=\frac{1}{|N(v)|}\sum_{u\in N(v)}s_{e},$ (37)
where the sum is over all edges incident to $v$, and $|N(v)|$ denotes the total number of neighbors. Edges where either function shows insufficient variation (below threshold) contribute zero rather than being excluded from the denominator.
The key distinction from previous measures lies in how sparse signal is handled. Return to the vertex with 30 neighbors where $y$ varies on all edges but $z$ shows meaningful signal on only one edge. Setting thresholds $\tau_{y}=\tau_{z}=0.05$ appropriate for the variation scales, the proportion-based measure yields $\text{cm}_{\text{prop}}(y,z)(v)=\pm 1/30\approx\pm 0.033$. This value correctly indicates that association is supported by sparse evidence: only $3\%$ of the neighborhood exhibits coordinated variation. As more edges develop signal, the coefficient magnitude increases proportionally, reaching $\pm 1$ only when directional concordance appears across the entire neighborhood. The fixed denominator prevents sparse signals from producing misleadingly large coefficients while maintaining interpretability as the proportion of edges showing directional agreement.
Threshold selection depends on data characteristics and inferential goals. For the response function $y$, thresholds can be set relative to overall variation: $\tau_{y}=c\cdot\text{sd}(y)$ for small $c$ (typically $0.05$ to $0.10$) filters changes smaller than $5-10\%$ of a standard deviation, removing numerical noise while preserving meaningful signal. For feature $z$ representing quantities like phylotype abundances that vary heterogeneously across features, adaptive per-feature thresholds based on quantiles of edge differences prove effective. Setting $\tau_{z_{j}}=Q_{0.25}(\{|\Delta_{e}z_{j}|:e\in E\})$ as the first quartile of absolute edge differences for feature $z_{j}$ filters the bottom $25\%$ of changes as noise while preserving variation in the upper three quartiles. This adaptive approach handles features with different baseline variability without requiring manual tuning. When thresholds are set to zero ($\tau_{y}=\tau_{z}=0$), the proportion-based measure reduces to the sign-based coefficient, recovering pure directional counting without magnitude filtering.
The proportion-based measure is particularly appropriate for binary or sparse features where the presence or absence of variation carries biological meaning. In microbiome studies, phylotypes may be absent from most samples but abundant where present, exhibiting extreme sparsity. The proportion-based coefficient with appropriate thresholds identifies regions where a phylotype’s presence correlates with outcome changes while appropriately penalizing associations supported by few samples. For continuous features exhibiting widespread variation, the correlation-type measure from Section 4.2 generally provides more sensitive detection of graded co-variation by weighting edges according to the magnitude of coordinated changes.
We summarize the comparison between normalization schemes. The correlation-type measure $\text{cm}_{\text{cor}}$ employs geometric mean normalization that provides stable coefficients across diverse signal patterns, interprets naturally as correlation between discrete derivatives, and weights edges by the magnitude of coordinated variation. It is recommended for general use with continuous features exhibiting heterogeneous but widespread variation. The proportion-based measure $\text{cm}_{\text{prop}}$ with thresholds explicitly filters weak signals, interprets as the proportion of neighborhood edges showing directional agreement, and naturally handles sparse or binary features where signal prevalence matters. It is recommended when explicit signal filtering is scientifically justified or when features exhibit extreme sparsity. The absolute value normalization $\text{cm}_{\text{abs}}$ becomes unstable with sparse signals and is not recommended for practical use. The sign-based normalization $\text{cm}_{\text{sign}}$ discards magnitude information that typically proves essential for distinguishing meaningful associations from noise and is not recommended for practical use. In applications where the choice is unclear, we suggest computing both $\text{cm}_{\text{cor}}$ and $\text{cm}_{\text{prop}}$ with conservative thresholds, examining their agreement to assess whether magnitude weighting versus prevalence counting yields substantively different conclusions.